# First principle prediction of structural distortions in the cuprates and their impact on the electronic structure

### v2

@@ Appendix A Three-band versus one-band model
Conventionally, the low-energy effective model can also be constructed from the Cu $d_{x^{2}-y^{2}}$ orbitals and the O $p_{\sigma}$ ($p_{x/y}$ orbitals pointing toward Cu atoms). This model is often known as the “three-band model” or Emery model Emery (1987). In this section, we will compare the three-band model with the one-band model in the main text. It turns out that they provide equivalent descriptions for the bands around the Fermi level. Hence, the choice of different effective models does not affect the results in the main text.
@@
Finally, we study the local moments and the electron occupancies in the three-band model as a comparison to the one-band model results in Fig. 5. Here we focus on the same stripe order shown in Fig. 5(b) and Wannierize it to the three-band model. Fig. 15(a) shows the local moments of different Cu $d_{x^{2}-y^{2}}$ and O $p_{\sigma}$ Wannier orbitals in the unit cell. The dominant local moments are on Cu, showing a consistent wavy pattern as the one-band model results in Fig. 5(c). Fig. 15(b) shows the electron occupancies of different Wannier orbitals, where both Cu and O Wannier orbitals show a consistent wavy pattern as the one-band model results in Fig. 5(d). This analysis provides a consistent picture in a prior work on YBCO Zhang et al. (2020).
@@ Appendix B Stripe order states and their energies in hole-doped system
As discussed in the main text, we have successfully identified many stripe-ordered states that are almost degenerate with the G-AFM state in our DFT calculations. We conclude that there is likely strong spin fluctuations in this overdoped system. Such spin fluctuations in cuprates are also detected by experiments Zaanen et al. (1996); Lyons et al. (1988); Lyons and Fleury (1988); Brom and Zaanen (2003); Zhu et al. (2023). In this section, we provide more details about these competing ordered states.
@@
In addition to the bond-centered stripe orders, a previous DFT study on YBCO has also identified site-centered stripe orders Zhang et al. (2020), where all domain walls are at Cu atomic sites. However, such site-centered stripe orders are unstable in Bi-2212 according to our calculations. Instead, we find several “hybrid” stripe orders, where domain walls appear both at the Cu-Cu bonds and the Cu atomic sites as illustrated in Fig. 16(g-i). The red dashed double arrows show the domain walls on Cu atomic sites, where the local moments are almost zero. Interestingly, we find several Cu sites with very small local moments marked by the red arrows, which are very common in the stripe patterns such as those in Fig. 16(c), (f), and (h). The physical origin of this phenomenon is likely due to the superlattice modulation distortion which modulates the local environments of the Cu atoms.
In principle, there are many more possible configurations of bond-centered stripe orders, because one can take combinations of the existing stripe patterns for the upper and lower CuO2 layers. We have tested 9 additional bond-centered patterns by taking these combinations and find energies of 1.65$\sim$3.40 meV/Cu above the G-AFM state. These stripe-order states, constructed by such combinations, generally show higher total energies than the existing stripe-order states (but all of them are very close in energy). All these calculations indicate the energetic tendency for existence of strong spin fluctuations in this system: as we show in the main text, one has to go beyond DFT to take account of the spin fluctuation effect to make quantitative predictions of ARPES spectra.

### v3

@@ Appendix A Three-band versus one-band model
Conventionally, the low-energy effective model can also be constructed from the Cu $d_{x^{2}-y^{2}}$ orbitals and the O $p_{\sigma}$ ($p_{x/y}$ orbitals pointing toward Cu atoms). This model is often known as the “three-band model” or Emery model [164]. In this section, we will compare the three-band model with the one-band model in the main text. It turns out that they provide equivalent descriptions for the bands around the Fermi level. Hence, the choice of different effective models does not affect the results in the main text.
@@
Finally, we study the local moments and the electron occupancies in the three-band model as a comparison to the one-band model results in Fig. 5. Here we focus on the same stripe order shown in Fig. 5(b) and Wannierize it to the three-band model. Fig. 16(a) shows the local moments of different Cu $d_{x^{2}-y^{2}}$ and O $p_{\sigma}$ Wannier orbitals in the unit cell. The dominant local moments are on Cu, showing a consistent wavy pattern as the one-band model results in Fig. 5(c). Fig. 16(b) shows the electron occupancies of different Wannier orbitals, where both Cu and O Wannier orbitals show a consistent wavy pattern as the one-band model results in Fig. 5(d). This analysis provides a consistent picture in a prior work on YBCO [96].
@@ Appendix B Stripe order states and their energies in hole-doped system
As discussed in the main text, we have successfully identified many stripe-ordered states that are almost degenerate with the G-AFM state in our DFT calculations. We conclude that there is likely strong spin fluctuations in this overdoped system. Such spin fluctuations in cuprates are also detected by experiments [141, 130, 133, 131, 132]. In this section, we provide more details about these competing ordered states.
@@
In addition to the bond-centered stripe orders, a previous DFT study on YBCO has also identified site-centered stripe orders [96], where all domain walls are at Cu atomic sites. However, such site-centered stripe orders are unstable in Bi-2212 according to our calculations. Instead, we find several “hybrid” stripe orders, where domain walls appear both at the Cu-Cu bonds and the Cu atomic sites as illustrated in Fig. 17(g-i). The red dashed double arrows show the domain walls on Cu atomic sites, where the local moments are almost zero. Interestingly, we find several Cu sites with very small local moments marked by the red arrows, which are very common in the stripe patterns such as those in Fig. 17(c), (f), and (h). The physical origin of this phenomenon is likely due to the superlattice modulation distortion which modulates the local environments of the Cu atoms.
In principle, there are many more possible configurations of bond-centered stripe orders, because one can take combinations of the existing stripe patterns for the upper and lower CuO2 layers. We have tested 9 additional bond-centered patterns by taking these combinations and found energies of 1.65$\sim$3.40 meV/Cu above the G-AFM state. These stripe-order states, constructed by such combinations, generally show higher total energies than the existing stripe-order states (but all of them are very close in energy). All these calculations indicate the energetic tendency for the existence of strong spin fluctuations in this system: as we show in the main text, one has to go beyond DFT to take account of the spin fluctuation effect to make quantitative predictions of ARPES spectra.